Subject: Minutes of the APO Users' Committee Meeting, March 10, 1997

From: elt@astro.Princeton.EDU

Submitted: Fri, 28 Mar 1997 10:23:18 -0500 (EST)

Message number: 135 (previous: 134, next: 136 up: Index)

Minutes of the 3.5-m User Committee meeting on March 10, 1997.

Present: Turner (chair), Walterbos (minutes), Hobbs, Stubbs, Gillespie, 
	 Secker, Uomoto

1. Instrument fixes and upgrades. 

New improvements management structure has been setup; 3-yr plan and urgent fixes
are two separate categories.

Stubbs is in charge of supervising activities for the 3-yr plan within
the established time table.

Klaene will have similar role for urgent repairs on the mountain.
Weekly teleconferences between parties involved (alternating between
3-yr plan and on-site fixes) will be used to manage and coordinate
this. APO mail exploders are good place to provide input to
this. People should still jump in and help, and volunteer for
that. You can send e-mail to find out on status of various issues to:

Users can also just call the site.

Also checking day and night logs is a good place to get
information. E.g. work on instruments will be logged in the day logs.

2. WSMR test last week was successful. Telescope tracked target
meeting their requirements. WSMR participation is helpful to
completion of 3-yr plan objectives.

3. SPICAM. Software redone, previous manual is obsolete. It is the
intent to make this a regularly scheduled instrument in Q2. Two
filterwheels will hold 6 filters each, 3"x3", adapter for 2"x2" is
being worked on. Thickness up to 12 mm or so OK. Filterset: UBVRI
procrured, permanently mounted in wheel 1. So other filters will be
installed in wheel 2. Sloan filter set still to be decided upon. No
narrow-band 3"x3" available, but several institutions have 2"X2"
narrow-bands. Alan has some moderate-band 3"x3" filters.

We need to make a full filter inventory for the 3.5-m.

4. Computer problems. Intermittent dropout in communications between
telescope control computers and others. Seems to have been fixed
now. Another problem: operations control computer system shows more
problems than before shutdown. Perhaps due to upgrade of operating
system on MC. Fowler is trying to find origin of problem.

5. Focus shifts: primary mirror's ventilation system changed since
shutdown. Also temperature shifts seem to cause problems; perhaps due
to new sensor system that uses sensors in different
places. Calibration used by software to adjust focus may need to be

6. Schedules for Q2. Scheduling is working less well, because number
of requests for specific slots is growing. There is often overlap
between different observers requesting for same slots. Lower ranked
proposals are therefore not very well scheduled, if at all. Possible
modification: request only 2/3 or 3/4 of the time and declaring rest
open time. For example, current Q2 schedule has 15 unscheduled half
nights due to unschedulable time.

Or have everyone give 2/3 request and then they receive back 1/3 which
they schedule themselves.

Discuss this with schedulers.

Engineering time: 3-yr plan activities may lead to bumped observers,
because things cannot be scheduled 3 months in advance. We ask for the
indulgence of the users, to make things better in optimum ways.

7. CHAOS update: laser is scheduled to be taken off after April
run. Prospects after that are unknown. Echelle testing on the testing
may start later this year, into fall quarter.

8. Dormitories: relax 5-day warning for reservations to 2-day
warning. Policy is: you take what you get if you give less than 2-days
(could be Sunspot or trailer accommodation).

9. Closing criteria: up to Observing specialist, not to the
observer. Harassing the telescope specialist is not allowable under
any circumstances. This includes not asking every 15 minutes about the
status. Let the telescope specialist know where you are, and they will
contact you. Specific written policy will be posted on Web.

10. Dust monitoring system in place now. Needs to be calibrated to
translate it into actual conditions required for shut down because of
dust. Issues include: how easy can certain dust particles be cleaned
off. Snow conditions lead to unavoidable closure, if snow is on the
roof. Are looking into upgrading roof heating system to melt snow off
as it falls. Dome is difficult for dealing with snow removal (flat
roof, telescope has fixed orientation with respect to slit).

11. Minutes of last meeting approved. Next meeting will be on

APO APO APO APO APO  Apache Point Observatory 3.5m  APO APO APO
APO  This is message 135 in the apo35-general archive. You can find
APO  the archive on
APO  To join/leave the list, send mail to
APO  To post a message, mail it to